OpenStudy (anonymous):

How many solutions does the equation 3x + 6 = −1 − 3 + 4x have?

OpenStudy (anonymous):

Please help

OpenStudy (hungergames13):

1 answer, -2

OpenStudy (anonymous):

really?

OpenStudy (hungergames13):

i belive so

OpenStudy (hungergames13):

I don't get anything else

OpenStudy (anonymous):

Wow that was fast!

OpenStudy (hungergames13):

:)

OpenStudy (anonymous):

What about ninjaslice?

OpenStudy (ninjaslice):

If you add me on skype I can tell you =) I can't on here.... elijah.math I got suspended for giving out free answers lol

OpenStudy (anonymous):

I got a 100!!!!

OpenStudy (hungergames13):

yeah:D

OpenStudy (anonymous):

Thank you so much! Hungergames

OpenStudy (anonymous):

13

OpenStudy (hungergames13):

your welcome

OpenStudy (anonymous):

do you know what bumping a question means?

OpenStudy (ninjaslice):

@hungergames13 Please do not give a solid answer, or you will be suspended sadly, I got suspended for doing the same thing. They call it DUmping

OpenStudy (ninjaslice):

But Yes, his answer is correct =)

OpenStudy (hungergames13):

bumping quistion means it bumps it to the top

OpenStudy (hungergames13):

question

OpenStudy (anonymous):

Ohhhhhhhh... that makes sense!

OpenStudy (hungergames13):

:)

OpenStudy (anonymous):

Can you help me with another question?

OpenStudy (hungergames13):

sure

OpenStudy (anonymous):

How do you insert a graph?

OpenStudy (hungergames13):

what do you mean?

OpenStudy (anonymous):

the problem involves a graph but i don't know how to show you.

OpenStudy (hungergames13):

ohh ok all you have to do is click on the graph but with the left side of your mouse and click on that one that says save image as.....

OpenStudy (hungergames13):

cick save and then come back here and go down to the bottom of were you type and click attach file

OpenStudy (anonymous):

The graph shows two lines, A and B: Based on the graph, which statement is correct about the solution to the system of equations for lines A and B? A: (0, 6) is the solution to both lines A and B. B: (0, 6) is the solution to line B but not to line A. C: (2, 2) is the solution to both lines A and B. D: (2, 2) is the solution to line A but not to line B.

OpenStudy (hungergames13):

hold on a sec

OpenStudy (anonymous):

ok

OpenStudy (hungergames13):

I think its A because...

OpenStudy (ninjaslice):

Watch out answering questions directly please, it is against terms of service.

OpenStudy (hungergames13):

I am trying to explain it but hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What a bout: A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 3y + 4z = 1 Which of the following is a possible step used in eliminating the y-term? (y + z = 6) ⋅ 4 (3y + 4z = 1) ⋅ 4 (y + z = 6) ⋅ −3 (3y + 4z = 1) ⋅ 3

OpenStudy (anonymous):

Sorry if I'm being too pushy.

OpenStudy (anonymous):

:(

OpenStudy (hungergames13):

no it's ok:)

OpenStudy (hungergames13):

one sec

OpenStudy (anonymous):

nevermind i found the answer

OpenStudy (anonymous):

could you help me with the next one?

OpenStudy (hungergames13):

ok this one is kinda hard:D

OpenStudy (hungergames13):

sure:)

OpenStudy (anonymous):

What a bout: A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 3y + 4z = 1 Which of the following is a possible step used in eliminating the y-term? (y + z = 6) ⋅ 4 (3y + 4z = 1) ⋅ 4 (y + z = 6) ⋅ −3 (3y + 4z = 1) ⋅ 3

OpenStudy (hungergames13):

this isn't the same one?

OpenStudy (anonymous):

ohh sorry wrong one

OpenStudy (hungergames13):

ohh ok xD

OpenStudy (anonymous):

A set of equations is given below: Equation E: a = 5b + 1 Equation F: a = 3b + 4 Which statement describes a step that can be used to find the solution to the set of equations? Equation F can be written as 5b + 1 = 3b + 4. Equation F can be written as b + 1 = 3b + 4. Equation F can be written as a = 3(a − 1) + 4. Equation F can be written as a = 5(a − 4) + 1.

OpenStudy (anonymous):

It is similar but not the same :)

OpenStudy (hungergames13):

ok

OpenStudy (hungergames13):

I think it's C

OpenStudy (anonymous):

Any thing? I'm not having luck :(

OpenStudy (anonymous):

oh ok

OpenStudy (hungergames13):

but I could be wrong

OpenStudy (hungergames13):

but i'm pretty sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have one more, but I promise you this is the last one ok. :)

OpenStudy (hungergames13):

it's ok

OpenStudy (anonymous):

Two lines, A and B, are represented by the equations given below: Line A: y = x − 6 Line B: y = 3x + 4 Which of the following shows the solution to the system of equations and explains why? (−5, −11), because the point satisfies both equations (−5, −11), because the point does not lie on any axis (−3, −5), because the point satisfies one of the equations (−3, −5), because the point lies between the two axes

OpenStudy (hungergames13):

I think it's A, yah

OpenStudy (anonymous):

Ok let me turn it il get back to you in a sec

OpenStudy (hungergames13):

ok hope you pass

OpenStudy (anonymous):

i got one wrong

OpenStudy (anonymous):

it was this one: A set of equations is given below: Equation E: a = 5b + 1 Equation F: a = 3b + 4 Which statement describes a step that can be used to find the solution to the set of equations? Equation F can be written as 5b + 1 = 3b + 4. Equation F can be written as b + 1 = 3b + 4. Equation F can be written as a = 3(a − 1) + 4. Equation F can be written as a = 5(a − 4) + 1.

OpenStudy (hungergames13):

ooohhhh sorry

OpenStudy (anonymous):

is okay i can retake it two more times

OpenStudy (hungergames13):

but dosn't that mean that you got an 80%?

OpenStudy (hungergames13):

or did you get another one wrong?

OpenStudy (anonymous):

no only one but there were only four questions so i got a 75

OpenStudy (hungergames13):

ohh ok

OpenStudy (hungergames13):

well hope you pass the next one

OpenStudy (anonymous):

ok thank you your help!!!!!!!!

OpenStudy (hungergames13):

your welcome! :D